OpenStudy (anonymous):

A 6-foot-tall man standing on the ground casts a shadow. The angle of the elevation of the Sun is 20°. How long is the shadow? Round your answer to the nearest tenth of a foot.

OpenStudy (anonymous):

If the ground is level, you will have a 90 degree or right triangle.

OpenStudy (anonymous):

\[x=\frac{6}{\tan(20)}\] use a calcular

OpenStudy (anonymous):

Is the scenario in the image I've attached correct?

OpenStudy (anonymous):

If so, then you can use trigonometry as satellite73 has shown.

OpenStudy (anonymous):

tan(20) will be equal to the adjacent side (the shadow length) divided by the opposite side. Rearranging gives the formula a=6/tan(20)

OpenStudy (anonymous):

\[6 \text{ Cot}\left[\frac{\pi }{9}\right] \]